### v1

@@ Abstract
Squeezed mechanical states are a valuable tool for quantum sensing and error correction in quantum computing, and a pivotal platform for tests of fundamental physics. Recently, solid state mechanical oscillators have been prepared in squeezed states using parametric interactions in both the microwave and optical regimes. It has long been predicted that a fast measurement rate comparable to the mechanical resonance frequency can prepare the oscillator under measurement into a quantum squeezed state. Despite decades of effort, this straightforward protocol is yet to be demonstrated in the quantum regime. Here, we use post-processing techniques to demonstrate preparation of a $50$ ng GaAs cantilever in a conditional classical squeezed state with a minimum uncertainty $(0.28\pm 0.18)$ dB above ($1.07\pm 0.04$ times) the zero point fluctuations, $3$ orders of magnitude closer to the quantum regime in variance than the previous record. This paves the way to real-time measurement-based preparation of macroscopic oscillators in quantum squeezed states, and can be adapted to mechanical systems as large as the kg-scale test masses of the Laser Interferometer Gravitational-Wave Observatory (LIGO).
@@
A quantum squeezed state is characterized by exhibiting one quadrature with uncertainty less than that of the zero point fluctuation (ZPF), at the expense of larger uncertainty in the conjugate quadrature. In the squeezed quadrature, measurement of a physical quantity is subjected to less noise than the corresponding unsqueezed ground state. Due to this noise reduction, for a mechanical system, the generation and observation of a quantum squeezed state is expected to provide a substantial resource for quantum sensing and quantum metrology [1, 2, 3], and for error correction in quantum computing [4, 5]. Furthermore, due to the relatively large mass of solid-state mechanical oscillators, mechanical quantum squeezed states are also a promising probe of the quantum-to-classical transition [6, 7, 8, 9], and test of quantumness of gravity [10].
@@
The key to preparing mechanical quantum squeezed states is to violate the equipartition of energy between the displacement $\hat{q}$ and the momentum $\hat{p}$ quadratures. Solid state mechanical oscillators have first been prepared in quantum squeezed states using reservoir engineering on microwave devices [11, 12, 13, 14]. By introducing coupling to superconducting qubits, non-Gaussian quantum squeezed states have been successfully prepared [15, 16]. In the optical domain, direct migration of the reservoir engineering technique to optical devices has achieved back-action evasion, but fell short of quantum squeezed state preparation due to excess noise such as optical heating [17]. Instead, inspired by ion trapping techniques [18], where quantum squeezed mechanical states are prepared by modulating the trapping potential, quantum squeezed center-of-mass motion of a particle levitated in laser fields was inferred after subtracting readout noise [19]. So far, all successful demonstrations of mechanical quantum squeezed state preparation can be classified as utilizing parametric interactions.
Another avenue for preparing mechanical quantum states is a continuous measurement-based preparation scheme, which has the advantage of simpler and steadier experimental control. This not only allows for a simplified experimental setup, but would also be directly applicable to complex instruments such as LIGO to avoid catastrophic instability. Feedback cooling of mechanical oscillators to the ground state using measurement rates faster than the thermal decoherence rate has already been achieved [20, 21, 22]. However, in these demonstrations, the equipartition of energy in the displacement and momentum quadratures is preserved due to the much faster mechanical oscillations compared to the measurement rates. In this regime, the rotating wave approximation is valid, precluding the preparation of squeezed states. It has been predicted more than two decades ago that a measurement rate $\Gamma_{\mathrm{meas}}$ comparable or faster than the resonance frequency of the mechanical oscillator $\Omega_{\mathrm{m}}$ can violate equipartition, and thus, allow for preparation of quantum squeezed states [23]. This approach has been refined over the last decades by incorporating state estimation and conditional state preparation [24, 25].
Measurement-Based Preparation of Mechanical Squeezed State
@@
For Gaussian states, the mean values of the estimators $\langle\hat{q}_{\mathrm{e}}\rangle$ and $\langle\hat{p}_{\mathrm{e}}\rangle$ and the (co)-variances of their corresponding errors $V_{\delta q\delta q}$, $V_{\delta p\delta p}$ and $C_{\delta q\delta p}$ uniquely define a state centered at the mean values in phase space with uncertainties given by the estimation errors [24, 25, 29]. As the estimators and estimation errors depend on the measurement record $\hat{y}$, the state is deemed a conditional state. The mean values directly inherit the stochastic nature of the measurement record $\hat{y}(t)$, resulting in a random walk of the conditional state center in phase space, within the distribution of a thermal state given by the thermal and back-action baths, as shown in Fig. 1(a). On the contrary, the variances of the estimation errors decrease deterministically in time, and converge to a steady state value, as more information is extracted from the measurement process [28]. The steady state value is determined by measurement rate $\Gamma_{\mathrm{meas}}$, thermal decoherence rate $\Gamma_{\mathrm{th}}$ and back-action rate $\Gamma_{\mathrm{ba}}$ shown in Fig. 1(b) and the mechanical frequency $\Omega_{\mathrm{m}}$ shown in Fig. 1(c). A conditional state prepared by causal filters (constructed using past information) can be transferred into an unconditional state through feedback in real time [30]. On the other hand, due to the requirement of future information, a conditional state prepared by non-causal filters is not accessible in real time, and cannot be transferred to an unconditional state by feedback. Moreover, non-causal filters may result in unphysical conditional states, which violates the Heisenberg uncertainty principle [31]. Therefore, for measurement-based preparation of physical mechanical states, estimate filters should be causal. In addition, to minimize the uncertainty, causal Wiener filters are the typical choices of $H_{q}$ and $H_{p}$. We follow this practice in our work.
@@
Figure 1: (a) Evolution of a conditional state. The axes are in dimensionless units defined as $\hat{Q}=\hat{q}/2q_{\mathrm{zpf}}$ and $\hat{P}=\hat{p}/2p_{\mathrm{zpf}}$, with $q_{\mathrm{zpf}}=\sqrt{\hbar/2m\Omega_{\mathrm{m}}}$ and $p_{\mathrm{zpf}}=\sqrt{\hbar m\Omega_{\mathrm{m}}/2}$, where $m$ is the effective mass and $\Omega_{\mathrm{m}}$ is the mechanical resonance frequency. The gray shade represents the estimation errors (i.e. the uncertainties of the conditional state), the green circle indicates the magnitude of the ZPF, and the dashed circle represents the thermal distribution. The center of the conditional state goes through a random walk, while the uncertainties reduce deterministically, converging to a steady state. (b) The uncertainties of the conditional state. Thermal decoherence rate $\Gamma_{\mathrm{th}}$ and back-action rate $\Gamma_{\mathrm{ba}}$ increases the uncertainty in the momentum quadrature, while the measurement rate $\Gamma_{\mathrm{meas}}$ decreases the uncertainty in the displacement quadrature. (c) The effect of quadrature rotation on the uncertainties of the conditional state.
@@
Now we discuss how the rates and mechanical frequency influence the steady state estimation errors. The Heisenberg-Langevin equations for a cavity optomechanical system are
@@
where $\hat{X}$ and $\hat{Y}$ are the amplitude and phase fluctuations of the intra-cavity optical field respectively; $\kappa$ is the angular cavity linewidth; $\Delta$ is the angular cavity detuning; $\hat{X}_{\mathrm{in}}$ and $\hat{Y}_{\mathrm{in}}$ are vacuum amplitude and phase noises, respectively, coupled into the intra-cavity field. Here, for simplicity, we do not distinguish the origin of the vacuum noises. For a complete model, please see Methods A. $G=\partial\Omega_{\mathrm{cav}}/\partial\hat{q}$ is the optomechanical interaction strength, with $\Omega_{\mathrm{cav}}$ the cavity angular resonance frequency; $\alpha=\sqrt{n_{\mathrm{cav}}}$ is the classical amplitude of the intra-cavity optical field, with $n_{\mathrm{cav}}$ the intra-cavity photon number; $m$ is the effective mass of the oscillator; $\Gamma_{\mathrm{m}}$ is the natural angular mechanical linewidth; and $\hat{\xi}_{\mathrm{th}}$ is the thermal force noise.
@@
The third term on the right-hand side of Eqn. 2 encodes the mechanical displacement $\hat{q}$ to the optical field. Extracting this information through optical detection refines the estimation of $\hat{q}$, therefore decreases the uncertainty $\delta\hat{q}$ at a rate of $\Gamma_{\mathrm{meas}}$, as shown in Fig. 1(b). $\Gamma_{\mathrm{meas}}$ is proportional to $n_{\mathrm{cav}}$, $\eta$, and $1/\kappa$, with $\eta$ the detection efficiency. At the same time, the second term on the right-hand side of Eq. 4 is the back-action noise, which directly couples only to the momentum quadrature at a rate of $\Gamma_{\mathrm{ba}}$. Similarly, the thermal force noise couples directly to the momentum quadrature at a rate of $\Gamma_{\mathrm{th}}=\Gamma_{\mathrm{m}}n_{\mathrm{th}}$ with $n_{\mathrm{th}}$ the thermal phonon occupancy of the environment. These two noises tend to increase uncertainty in $\delta\hat{p}$, as shown in Fig. 1(b). The three rates have the tendency of squeezing the mechanical state naturally. However, the quadrature rotation represented by Eq. 3 indirectly couples the back-action noise and the thermal force noise to $\hat{q}$, and competes with the process of uncertainty reduction through measurement, as illustrated in Fig. 1(c). For most optomechanical systems, $\Gamma_{\mathrm{meas}}\ll\Omega_{\mathrm{m}}$, the quadrature rotation averages away the squeezing from measurement, leaving a circular (non-squeezed) conditional state, as shown in the upper branch of Fig. 1(c). If the measurement rate is comparable to or faster than the angular frequency of the mechanical oscillator, i.e. $\Gamma_{\mathrm{meas}}\gtrsim\Omega_{\mathrm{m}}$, back-action and thermal noises coupled to $\hat{p}$ are rotated to $\hat{q}$ slower than the reduction of $\delta\hat{q}$ by measurement, and equipartition can be violated. Thus the squeezing from measurement is preserved, as shown in the lower branch of Fig. 1(c).
@@
The condition of $\Gamma_{\mathrm{meas}}\gtrsim\Omega_{\mathrm{m}}$ is easiest to achieve with low frequency mechanical oscillators. Due to excess classical noises, previous experimental efforts with low frequency mechanical oscillators were unable to achieve the preparation of quantum squeezed states, instead preparing states deemed classically squeezed, that is a state violating the equipartition between $\hat{q}$ and $\hat{p}$ but with no quadrature uncertainty below the ZPF [32, 33]. The previous record was set by [32], where the standard deviation of the displacement quadrature was 36 times of ZPF, i.e. 1296 times in variance.
@@
In this work, we demonstrate the preparation of a $50$ ng oscillator into a conditional mechanical classically squeezed state in post processing using casual Wiener filters for state estimation, with the lowest quadrature uncertainty in variance only $(0.28\pm 0.18)$ dB above ($1.07\pm 0.04$ times) the ZPF, where the uncertainty is one standard deviation of the variance. The conjugate quadrature uncertainty is $(8.98\pm 0.11)$ dB above ($7.9\pm 0.2$ times of) the ZPF, clearly demonstrating the violation of equipartition. This result is at the cutting edge of squeezing of the motional state of a mechanical oscillator via measurement, improving on previous record by 3 orders of magnitude. This is a significant milestone towards the quantum regime. Our work shines light on the problem of quantum-to-classical transition, especially the interplay between gravity and quantum states [6, 7, 10]. Due to the relatively large mass of our oscillator, it is also an ideal platform to demonstrate quantum enhanced accelerometry [34]. Because of the simple and steady control, our work also paves the way to preparing LIGO test masses into quantum squeezed states without the risk of catastrophic instability.
@@
Figure 2: (a) A simplified optical layout of the experiment. Input light is blue-detuned from the cavity resonance, stiffening the mechanical oscillator, and the system dynamics is stabilized by feeding back to the amplitude of the input light. Displacement information is readout by a balanced homodyne receiver in reflection. The photo current is saved for post processing. (b) The optomechanical resonator as imaged with a confocal laser microscope.
@@
The optomechanical resonator used in our experiment is a high reflectivity Bragg mirror formed of 23 periods of quarter-wave optical thickness GaAs and AlGaAs. The mirror consists of a circular pad with a nominal diameter of $70~{}\mu$m at the end of a nominally $55~{}\mu$m long, $8~{}\mu$m wide, and $220$ nm thick cantilever that is attached to a ledge of the GaAs substrate. The fundamental resonance of the mirror is $\Omega_{m}/2\pi=876$ Hz with a Q factor of $16,000$ and modal mass of $50$ ng. For more detail on the fabrication of the optomechanical resonator, see Refs [35, 36, 37, 38].
@@
The optomechanical resonator is cryogenically cooled to $50$ K and has a transmissivity of $450$ ppm at optical wavelength $\lambda=1064$ nm. This differs from the room temperature transmissivity of $250$ ppm in [39] due to thermorefractive effects in the Bragg mirror. The optomechanical resonator along with a standard highly reflecting mirror form a 1 cm long Fabry-Perot cavity that is placed under vacuum (~$10^{-8}$ Torr). The cavity has a linewidth of $\kappa/2\pi=1.77$ MHz. $79\%$ of the input light is spatially mode-matched to the cavity, and builds up to $0.25$ W of circulating power. The rest of the input light does not match the spatial mode of the cavity, and is promptly reflected. The influence of this mode mismatched part of light is considered in the model (see Methods A). The frequency of the input light is blue-detuned from the cavity resonance by a fractional detuning of $\delta_{\mathrm{cav}}=2\Delta/\kappa=0.24$. Blue-detuning creates a recoil radiation pressure force on the cantilever (i.e. an optical spring), which self-locks the cavity detuning [40]. To stabilize the mechanical resonator due to the blue-detuning, we apply feedback to an amplitude modulator before the cavity using the light exiting from the 1-cm mirror, as shown in Fig. 2. The optical spring and feedback trap the mechanical oscillator at a frequency of $\Omega_{\mathrm{eff}}/2\pi=113.7$ kHz. Optically stiffening the mechanical oscillator in this way suppresses thermal noise [41] (see Methods F), and pushes the resonance above the frequency band where low frequency technical noises dominate.
@@
Light exiting the cantilever is incident on a balanced homodyne receiver for phase quadrature detection. Photo current from the homodyne receiver is digitized for analysis at a 2 MHz sampling rate using a Red Pitaya STEMlab 125-14. The detection efficiency $\eta$ from the cantilever to photo current is $0.42$ due to losses in optics, imperfect homodyne visibility, and the quantum efficiency of the photodiodes. The cavity efficiency, intracavity power and detection chain give a high measurement rate of $61$ kHz, which is $54\%$ of the mechanical frequency, enough to violate the equipartition of energy in the displacement and momentum quadratures.
@@
The digitized homodyne photo current are saved as a time series. We fast Fourier transform the time series to get a voltage spectrum $S_{vv}[\Omega]$, which we subsequently calibrate from voltage to displacement $S_{yy}[\Omega]$ by matching the shot noise to the theoretical model (see Methods C). We then clean the calibrated spectrum by comparing it with auxiliary spectra measured at different cavity detunings (see Methods D), permitting the removal of noise common to all the spectra. As shown in Fig. 3(a), the cleaned displacement spectrum matches the model $S_{yy}[\Omega]$ well, excepting the parts influenced by non-stationary noise (e.g. to the left of the mechanical resonance). To avoid the influence of the non-stationary noise, we use the modeled $S_{yy}[\Omega]$ to compute the causal Wiener filters and to calculate the conditional state. We calculate the causal Wiener filters for displacement $\hat{q}$ using the Wiener-Hopf Technique [26] as
@@
where $M_{y}[\Omega]$ is the causal spectral factor of $S_{yy}[\Omega]$ and $M_{y}[\Omega]^{*}$ is the anti-causal spectral factor. $M_{y}[\Omega]M_{y}[\Omega]^{*}=S_{yy}[\Omega]$. $[...]_{+}$ takes the causal decomposition of the function in the bracket (see Chapter 3 of [26] for details). The cross spectra $S_{qy}[\Omega]$ come from the model. $H_{q}[\Omega]$ is shown in Fig. 3(b).
We then apply the causal Wiener filter to $S_{yy}$ by multiplying the two in the frequency domain to calculate the conditional displacement spectrum of the estimation error

### v2

@@ Abstract
Squeezed mechanical states are a highly-coveted resource for quantum-enhanced sensing and serve as a compelling platform for probing the interplay between gravity and quantum mechanics. It has been predicted that a mechanical oscillator can be prepared into a quantum squeezed state if the applied measurement rate is fast relative to its mechanical resonance frequency. However, the experimental feasibility of this protocol has remained uncertain because of the difficulty in achieving low-frequency oscillators with sufficiently strong read-out. Here, we demonstrate that a careful selection of parameters in an optomechanical system, combined with optimal filtering techniques, enables the preparation of a $50$ ng GaAs cantilever in a conditional classical squeezed state, achieving a minimum uncertainty of just $1.07\pm 0.04$ times the zero-point fluctuation level. This minimum variance is $3$ orders of magnitude smaller than what has been achieved in previous experiments using the same protocol. Although we do not fully achieve the quantum squeezed regime, our demonstration provides definitive evidence that a measurement-based protocol is a practical and effective approach for the real-time preparation of macroscopic oscillators in quantum squeezed states.
@@
Preparing and controlling quantum states in mechanical systems at mesoscopic and macroscopic scales has become a frontier in exploring the nature of the quantum-to-classical transition, which seeks to understand how classical behavior emerges from quantum laws [1, 2, 3, 4]. Investigating this transition involves probing various hypothetical decoherence mechanisms proportional to the mass of the quantum state. Moreover, preparing and stabilizing mesoscopic to macroscopic systems in quantum states provide a potential testbed for studying the quantum nature of gravity, where macroscopic superpositions and entanglement generated by gravity could offer insights into whether gravity itself is quantum [5, 6]. Recently, schemes employing Gaussian states have been proposed as promising alternatives with reduced experimental demands.[7, 8]. Mechanical oscillators in their quantum ground states - where fluctuations in displacement and momentum quadratures saturate the Heisenberg uncertainty relation - are a fundamental goal in the field of quantum state preparation [9]. In the ground state, energy is equally distributed between displacement and momentum quadratures, referred to as equipartition of energy.
Beyond this equipartitioned state lies the quantum squeezed state, where noise in one quadrature can be reduced below the zero-point fluctuation (ZPF) at the cost of increasing noise in the conjugate quadrature. As a result, squeezed states exhibit significantly greater sensitivity to changes in the noise-reduced quadrature than the ground state. This enhanced sensitivity makes them valuable for technologies that require ultraprecise measurements. [10, 9, 11]
@@
The key to preparing mechanical quantum squeezed states is to break the equipartition of energy between the displacement $\hat{q}$ and the momentum $\hat{p}$ quadratures. To date, solid-state mechanical oscillators have been prepared in quantum squeezed states using reservoir engineering on microwave devices [12, 13, 14, 15]. Furthermore, non-Gaussian quantum squeezed states have been successfully prepared by introducing coupling to superconducting qubits [16, 17]. In the optical domain, the approach analogous to reservoir engineering has successfully achieved back-action evasion but falls short by several orders of magnitude in the achievable squeezing due to excess noise, such as optical heating [18]. In contrast, inspired by ion trapping techniques [19], where quantum-squeezed mechanical states are prepared by modulating the trapping potential, quantum squeezing of the center-of-mass motion of a particle levitated in laser fields has been both inferred (after readout noise subtraction [20]) and directly measured [21]. So far, all successful demonstrations of mechanical quantum-squeezed state preparation have relied on parametric interactions.
Another avenue for preparing mechanical quantum states is continuous measurement, which offers simpler and steadier experimental control.
Feedback cooling of mechanical oscillators to the ground state using measurement rates faster than the thermal decoherence rate has already been achieved [22, 23, 24]. However, in these demonstrations, energy equipartition between the displacement and momentum quadratures is maintained because mechanical oscillations are much faster than the measurement rates, preventing the emergence of observable squeezing. In this regime, the rotating wave approximation is valid, implying that the dynamics do not support the preparation of squeezed states.
More than two decades ago, it was predicted that a measurement rate $\Gamma_{\mathrm{meas}}$ comparable to or exceeding the mechanical resonance frequency $\Omega_{\mathrm{m}}$ can violate energy equipartition and enable the preparation of quantum-squeezed states [25]. This approach has been refined over the last decades by incorporating state estimation and conditional state preparation [26, 27].
Measurement-Based Preparation of a Squeezed Mechanical State
State estimation is a widely used technique for extracting physical observables from noisy measurements and constitutes a central element in this work. A more detailed introduction can be found in Methods B. For any linear continuous measurement of displacement $\hat{q}$, the measurement outcome can be normalized to displacement units as $\hat{y}(t)=\hat{q}(t)+\hat{z}(t)$, where $\hat{z}(t)$ is the normalized measurement imprecision noise. To extract $\hat{q}(t)$ from a noisy measurement record $\hat{y}(t)$, an estimation filter $H_{q}(t)$ can be constructed and applied using a priori knowledge of the system, such as system parameters and noise spectra [28, 29, 30]. The displacement estimator is subsequently given by $\hat{q}_{\mathrm{e}}(t)=H_{q}*\hat{y}(t)$, with * denoting convolution. The error of the estimate is given by $\delta\hat{q}(t)=\hat{q}(t)-\hat{q}_{\mathrm{e}}(t)$. One can similarly define $\hat{p}_{\mathrm{e}}$ and $\delta\hat{p}$ for momentum $\hat{p}$.
@@
For Gaussian states, the mean values of the estimators, $\langle\hat{q}_{\mathrm{e}}\rangle$ and $\langle\hat{p}_{\mathrm{e}}\rangle$, along with the (co)variances of their corresponding errors, $V_{\delta q\delta q}$, $V_{\delta p\delta p}$, and $C_{\delta q\delta p}$, uniquely define a state centered at the estimated mean values in phase space, with uncertainties determined by the estimation errors [26, 27, 31]. Since estimators and their associated errors depend on the measurement record $\hat{y}$, the resulting state is termed a conditional state. The center of this conditional state undergoes a random walk in phase space, distributed according to a thermal state defined by thermal and back-action baths, as illustrated in Fig. 1(a). In contrast, the variances of the estimation errors decrease deterministically over time and converge to steady-state values as more information is acquired through measurement [30]. A conditional state prepared using causal filters, which rely solely on past information, can be converted into an unconditional state through real-time feedback [32]. Consequently, causal conditional states are typically employed for state preparation.
@@
Figure 1: (a) Evolution of a conditional state. The axes use dimensionless units, defined as $\hat{Q}=\hat{q}/q_{\mathrm{zpf}}$ and $\hat{P}=\hat{p}/p_{\mathrm{zpf}}$, where the zero-point fluctuations are given by $q_{\mathrm{zpf}}=\sqrt{\hbar/2m\Omega_{\mathrm{m}}}$ and $p_{\mathrm{zpf}}=\sqrt{\hbar m\Omega_{\mathrm{m}}/2}$. The gray shaded area represents the estimation errors (i.e. the uncertainties of the conditional state), the green disk marks the magnitude of the ZPF, and the dashed circle represents the thermal distribution. The center of the conditional state undergoes a random walk, while the uncertainties decrease deterministically, converging to a steady state. (b) Uncertainties of the conditional state. The thermal decoherence rate $\Gamma_{\mathrm{th}}$ and back-action rate $\Gamma_{\mathrm{ba}}$ increase the uncertainty in the momentum quadrature, while the measurement rate $\Gamma_{\mathrm{meas}}$ decreases the uncertainty in the displacement quadrature. (c) The effect of quadrature rotation on the uncertainties of the conditional state.
@@
In the following, we discuss how the rates and mechanical frequency influence the steady-state estimation errors. The Heisenberg-Langevin equations for a cavity optomechanical system are
@@
where $\hat{X}$ and $\hat{Y}$ are the amplitude and phase fluctuations of the intra-cavity optical field respectively; $\kappa$ is the angular cavity linewidth; $\Delta$ is the angular cavity detuning; $\hat{X}_{\mathrm{in}}$ and $\hat{Y}_{\mathrm{in}}$ are vacuum amplitude and phase noises, respectively, coupled into the intra-cavity field. For simplicity, we do not distinguish between the different sources of vacuum noise. For a complete model, please see Methods A. $G=\partial\Omega_{\mathrm{cav}}/\partial\hat{q}$ is the optomechanical interaction strength, with $\Omega_{\mathrm{cav}}$ the cavity angular resonance frequency; $\alpha=\sqrt{n_{\mathrm{cav}}}$ is the classical amplitude of the intra-cavity optical field, with $n_{\mathrm{cav}}$ the intra-cavity photon number; $m$ is the effective mass of the oscillator; $\Gamma_{\mathrm{m}}$ is the natural angular mechanical linewidth; and $\hat{\xi}_{\mathrm{th}}$ is the noise due to the thermal force.
@@
The third term on the right-hand side of Eqn. 2 encodes the mechanical displacement $\hat{q}$ into the optical field. Extracting this information via optical detection improves the estimation of $\hat{q}$ and thereby decreases the uncertainty $\delta\hat{q}$ at a rate of $\Gamma_{\mathrm{meas}}$, as shown in Fig. 1(b). $\Gamma_{\mathrm{meas}}$ is proportional to $n_{\mathrm{cav}}$, $\eta$, and $1/\kappa$, with $\eta$ the detection efficiency. At the same time, the second term on the right-hand side of Eq. 4 is the back-action noise, which directly couples only to the momentum quadrature at a rate of $\Gamma_{\mathrm{ba}}$. Similarly, the thermal force noise couples directly to the momentum quadrature at a rate of $\Gamma_{\mathrm{th}}=\Gamma_{\mathrm{m}}n_{\mathrm{th}}$ with $n_{\mathrm{th}}$ the thermal phonon occupancy of the environment. These two noises tend to increase the uncertainty in $\delta\hat{p}$, as shown in Fig. 1(b). $\Gamma_{\mathrm{meas}}$, $\Gamma_{\mathrm{ba}}$ and $\Gamma_{\mathrm{th}}$ naturally tend to squeeze the mechanical state. However, the quadrature rotation described by Eq. 3 indirectly couples the back-action noise and thermal force noise to $\hat{q}$, counteracting the uncertainty reduction achieved through measurement, as illustrated in Fig. 1(c). In most optomechanical systems, where $\Gamma_{\mathrm{meas}}\ll\Omega_{\mathrm{m}}$, the quadrature rotation averages out the squeezing induced by measurement, resulting in a circular (non-squeezed) conditional state, as shown in the upper branch of Fig. 1(c). If the measurement rate is comparable to or faster than the angular frequency of the mechanical oscillator, i.e. $\Gamma_{\mathrm{meas}}\gtrsim\Omega_{\mathrm{m}}$, back-action and thermal noises coupled to $\hat{p}$ are rotated to $\hat{q}$ slower than the reduction of $\delta\hat{q}$ by measurement, and equipartition can be violated. Thus the squeezing from measurement is preserved, as shown in the lower branch of Fig. 1(c).
@@
The condition $\Gamma_{\mathrm{meas}}\gtrsim\Omega_{\mathrm{m}}$ is most readily achieved with low-frequency mechanical oscillators. However, due to excess classical noise, previous experiments with such oscillators have been unable to prepare quantum-squeezed states. Instead, they produced states considered classically squeezed states that violate energy equipartition between $\hat{q}$ and $\hat{p}$ but do not exhibit quadrature uncertainties below the zero-point fluctuations (ZPF) [33, 34]. Along the coveted path of quantum-squeezed motional states, the previous record set by [33] achieved a standard deviation of the displacement quadrature that was 36 times ZPF, i.e. 1296 times the variance.
@@
In this work, we demonstrate the preparation of a $50$ ng oscillator in a classically squeezed conditional mechanical state using causal Wiener filters in subsequent analysis. (For a linear system driven by stationary Gaussian noise, it is well known that the Wiener filter minimizes the mean square of estimation errors [28].) The lowest achieved quadrature uncertainty in variance was merely $(0.28\pm 0.18)$ dB above the ZPF, corresponding to a factor of $(1.07\pm 0.04)$ times the ZPF, where the uncertainty reflects one standard deviation of the variance. The uncertainty in the conjugate quadrature was $(8.98\pm 0.11)$ dB above the ZPF, or $(7.9\pm 0.2)$ times of the ZPF, clearly demonstrating a violation of energy equipartition. This result represents a state-of-the-art achievement in measurement-based motional state squeezing of a mechanical oscillator, surpassing the previous record by three orders of magnitude. It marks a significant milestone toward reaching the quantum regime. Our work shines light on the problem of the quantum-to-classical transition, especially the interplay between gravity and quantum states, by providing a heavy and easy-to-use sensor [1, 2, 7]. Due to the relatively large mass of our oscillator, it is also an ideal platform to demonstrate quantum-enhanced accelerometry [35]. Moreover, complex optomechanical systems like LIGO are often plagued by instabilities that must be controlled. Our system includes a relatively simple control system, paving the way to preparing LIGO test masses in quantum-squeezed states without the risk of catastrophic instability.
@@
Figure 2: (a) Simplified optical layout of the experiment. AM, PID, $H_{q}$ and $H_{p}$ stand for optical amplitude modulator, PID servo, causal Wiener filter for $\hat{q}$ and $\hat{p}$ respectively. The input light is blue-detuned from the cavity resonance, stiffening the mechanical oscillator. The system dynamics are stabilized by feeding back to the amplitude of the input light. Displacement information is read out by a balanced homodyne receiver in reflection. The photocurrent is saved for subsequent analysis. (b) The optomechanical resonator as imaged with a microscope.
@@
The optomechanical resonator used in our experiment is a high-reflectivity Bragg mirror formed of 23 periods of quarter-wave optical thickness GaAs and AlGaAs. The mirror consists of a circular pad with a nominal diameter of $70~{}\mu$m at the end of a nominally $55~{}\mu$m long, $8~{}\mu$m wide, and $220$ nm thick cantilever that is attached to a ledge of the GaAs substrate. The fundamental resonance of the mirror is $\Omega_{m}/2\pi=876$ Hz with a Q-factor of $16,000$ and modal mass of $50$ ng. For more detail on the fabrication of the optomechanical resonator, see Refs [36, 37, 38, 39].
@@
The optomechanical resonator is cryogenically cooled to $50$ K and has a transmissivity of $450$ ppm at optical wavelength $\lambda=1064$ nm. This differs from the room temperature transmissivity of $250$ ppm in [40] due to thermorefractive effects in the Bragg mirror. The optomechanical resonator, together with a standard highly reflective mirror, forms a 1-cm-long Fabry-Perot cavity that is placed under vacuum at ~$10^{-8}$ Torr. The cavity has a linewidth of $\kappa/2\pi=1.77$ MHz. $79\%$ of the input light is spatially mode-matched to the cavity, and builds up to $0.25$ W of circulating power. The rest of the input light does not match the spatial mode of the cavity, and is promptly reflected. The influence of this mode mismatched part of light is considered in the model (see Methods A). The frequency of the input light is blue-detuned from the cavity resonance by a fractional detuning of $\delta_{\mathrm{cav}}=2\Delta/\kappa=0.24$. Blue-detuning creates a recoil radiation pressure force on the cantilever (i.e. an optical spring), which self-locks the cavity detuning [41]. To stabilize the mechanical resonator due to the blue-detuning, we apply feedback to an amplitude modulator before the cavity using the light exiting from the 1-cm mirror, as shown in Fig. 2. The optical spring and feedback trap the mechanical oscillator at a frequency of $\Omega_{\mathrm{eff}}/2\pi=113.7$ kHz. Optically stiffening the mechanical oscillator in this way suppresses thermal noise [42] (see Methods G), and pushes the resonance above the frequency band where low frequency technical noises dominate.
@@
Light exiting the cantilever is directed to a balanced homodyne receiver for phase quadrature detection. The resulting photocurrent is digitized for analysis at a 2 MHz sampling rate using a Red Pitaya STEMlab 125-14 for subsequent analysis. The overall detection efficiency, $\eta=0.42$, accounts for optical losses, imperfect homodyne visibility, and the finite quantum efficiency of the photodiodes. The combined effects of cavity efficiency, intracavity power, and detection chain yield a high measurement rate of 61 kHz, corresponding to 54% of the mechanical resonance frequency. This rate is sufficient to violate energy equipartition between the displacement and momentum quadratures.
@@
The digitized homodyne photocurrent is saved as a time series. We apply a fast Fourier transform to the time series to obtain the voltage spectrum $S_{vv}[\Omega]$, which is then calibrated to the displacement spectrum $S_{yy}[\Omega]$ by matching the shot noise to the theoretical model (see Methods D). We further refine the calibrated spectrum by comparing it with auxiliary spectra measured at different cavity detunings (see Methods E), allowing for the removal of noise common to all spectra. As shown in Fig. 3(a), the cleaned displacement spectrum matches the model $S_{yy}[\Omega]$ well, except in regions affected by non-stationary noise (e.g. frequencies slightly below the mechanical resonance). To mitigate the influence of non-stationary noise, we use the modeled $S_{yy}[\Omega]$ to compute the causal Wiener filters and to calculate the conditional state. We calculate the causal Wiener filters for displacement $\hat{q}$ using the Wiener-Hopf Technique [28] as
@@
where $M_{y}[\Omega]$ is the causal spectral factor of $S_{yy}[\Omega]$ and $M_{y}[\Omega]^{*}$ is the anti-causal spectral factor, such that $M_{y}[\Omega]M_{y}[\Omega]^{*}=S_{yy}[\Omega]$. $[...]_{+}$ takes the causal decomposition of the function in the bracket (see Chapter 3 of [28] for details). The cross spectrum $S_{qy}[\Omega]$ is derived directly from the theoretical model. $H_{q}[\Omega]$ is shown in Fig. 3(b).
We then apply the causal Wiener filter to $S_{yy}$ by multiplying them in the frequency domain, yielding the conditional displacement spectrum of the estimation error